# Nonlinear MPC for Feedback-Interconnected Systems: a Suboptimal and Reduced-Order Model Approach

### v1

@@ I INTRODUCTION
The main drawback of MPC schemes lies in their computational complexity, which stems from the need to solve an optimal control problem online at each sampling instant. Consequently, a central challenge in current MPC research is to mitigate this limitation. In this context, we focus on two strategies that are widely adopted in both theoretical studies and practical implementations: (i) model order reduction, and (ii) suboptimal control approaches. When dealing with complex systems characterized by hierarchical structures or multiple interconnected subsystems, the computational complexity of MPC can become prohibitive. In such cases, reduced-order models can be employed to simplify the optimization problem while still capturing the essential dynamics of the system. In this spirit, the works [4, 5, 6] propose multi-rate control architectures, in which high-level routines based on MPC operate on reduced-order models and provide setpoints to low-level controllers acting on the full-order ones. This paradigm, which is very common in robotic applications, has been formalized in the tutorial work [7]. In [8], the authors guarantee the stability of the full closed-loop system despite the use of a reduced model in the MPC optimizer. A reduced-order MPC scheme is presented in [9], where stability and constraint satisfaction are proven while explicitly accounting for model reduction errors, whose bounds can be computed, e.g., as in [10]. In [11], Galerkin projection methods are employed to operate on a low-dimensional surrogate model. The authors of [12] present a robust MPC framework for two-time-scale linear systems, where only a reduced-order model neglecting the fast dynamics is used in the MPC optimizer.
@@ II SCENARIO DESCRIPTION
where $x_{{t}}\in\mathbf{X}\subseteq\mathbb{R}^{n}$ is the target state, $\xi_{{t}}\in\boldsymbol{\Xi}\subseteq\mathbb{R}^{p}$ is the extra dynamics state, $f:\mathbf{X}\times\boldsymbol{\Xi}\times\mathbf{U}\times\mathbb{R}_{+}\to\mathbf{X}$ and $g:\boldsymbol{\Xi}\times\mathbf{X}\times\mathbf{U}\times\mathbb{R}_{+}\to\boldsymbol{\Xi}$ denote the target and extra dynamics, respectively, $u_{{t}}\in\mathbf{U}\subseteq\mathbb{R}^{m}$ is the control input, and $\delta>0$ is a parameter that allows to arbitrarily modulate the relative speed between the target state $x_{{t}}$ and the extra dynamics state $\xi_{{t}}$. We formally characterize this property as follows.
@@
There exists $\bar{\delta}_{1},L_{f}>0$ such that, for all $\delta\in(0,\bar{\delta}_{1})$, it holds
@@
$\left\|g(\xi,x,u,\delta)-g(\xi^{\prime},x^{\prime},u^{\prime},\delta)\right\|$ $\leq L_{g}\left\|\xi\!-\!\xi^{\prime}\right\|+L_{g}\left\|x\!-\!x^{\prime}\right\|$
$\hskip 11.38092pt+L_{g}\left\|u\!-\!u^{\prime}\right\|,$
for all $\xi,\xi^{\prime}\in\boldsymbol{\Xi}$, $x,x^{\prime}\in\mathbf{X}$, $u,u^{\prime}\in\mathbf{U}$.
@@
for all $x\in\mathbf{X}$, $u\in\mathbf{U}$, and $\delta>0$. Moreover, there exists a continuous function $U:\boldsymbol{\Xi}\to\mathbb{R}$ and $\bar{\delta}_{2}>0$ such that, for all $\delta\in(0,\bar{\delta}_{2})$, it holds
$a_{1}\|\tilde{\xi}\|^{2}\leq U(\tilde{\xi})\leq a_{2}\|\tilde{\xi}\|^{2}$ (3a)
$U\!\big(g(\tilde{\xi}\!+\!\xi_{\text{eq}}(x,u),x,u,\delta)\!-\!\xi_{\text{eq}}(x,u)\!\big)\!-\!U(\tilde{\xi})\!\!\leq\!\!-\!a_{3}\|\tilde{\xi}\|^{2}\!$ (3b)
@@
for all $\tilde{\xi},\tilde{\xi}^{\prime}\in\boldsymbol{\Xi}$, $x\in\mathbf{X}$, $u\in\mathbf{U}$ and some constants $a_{1},a_{2},a_{3},a_{4}>0$.
@@
By adopting a Forward-Euler discretization with sampling time $\delta>0$ and choosing the armature inductance scaled as $L=\tilde{L}\delta$ with $\tilde{L}>0$, we get a discretized counterpart of (5) described as
@@
Indeed, by observing (7a), we can see that system (6) satisfies Assumption 1 with $L_{f}=K_{t}/J$ and $L_{g}=\max\{1-R/\tilde{L},K_{e}/\tilde{L},1/\tilde{L}\}$. Moreover, by observing (7b), we also note that subsystem (6b) admits an equilibrium function $\xi_{\text{eq}}:\mathbb{S}^{1}\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ defined as
@@
III SUBOPTIMAL AND REDUCED-ORDER MPC: STRATEGY DESCRIPTION AND STABILITY GUARANTEES
@@ III-A MPC Strategy Description
The control input $u_{{t}}$ is iteratively computed through a suboptimal MPC scheme applied to the reduced system, thereby adhering to and further reinforcing the rationale of computational efficiency motivating the use of $f_{R}$. In detail, given a time horizon $T\in\mathbb{N}$, at each iteration ${t}\in\mathbb{N}$, we consider the finite-horizon optimal control problem
$\begin{aligned} \min_{\{u_{{\tau}}\}_{{\tau}=0}^{T-1}}\quad&\sum_{{\tau}=0}^{T-1}\ell\left(x_{{\tau}},u_{{\tau}}\right)+\ell_{T}\left(x_{T}\right)\\ \text{s.t.}\quad&x_{0}=x_{{t}}\\ &x_{{{\tau}+1}}=f_{R}(x_{{\tau}},u_{{\tau}},\delta),\quad{\tau}=0,\ldots,T-1\\ &u_{{\tau}}\in\mathcal{U},\quad x_{{{\tau}+1}}\in\mathcal{X},\quad{\tau}=0,\ldots,T-1\\ &x_{T}\in\mathcal{X}_{f},\end{aligned}$ (10)
where $\ell:\mathbf{X}\times\mathbf{U}\to\mathbb{R}_{+}$ is the stage cost, $\ell_{T}:\mathbf{X}\to\mathbb{R}_{+}$ is the terminal cost, $\mathcal{U}\subseteq\mathbf{U}$ is the input constraint set, $\mathcal{X}\subseteq\mathbf{X}$ is the state constraint set, and $\mathcal{X}_{f}\subseteq\mathbf{X}$ is the terminal constraint set. Problem (10) is addressed suboptimally. More in detail, we consider an optimizer dynamics of the form
$z_{{{t}+1}}$ $=\mathcal{T}(z_{{t}},x_{{t}})$ (11a)
@@
where $z_{{t}}\in\mathbf{Z}\subseteq\mathbb{R}^{Tm}$ is the optimizer state, $\mathcal{T}:\mathbf{Z}\times\mathbf{X}\to\mathbf{Z}$ is the optimization algorithm, and $T\in\mathbb{N}$ is the prediction horizon. As customary in MPC schemes, only the first element of the $z_{{t}}$ sequence is applied to the plant as control input, hence the projection map $\Pi:\mathbf{Z}\to\mathbf{U}$ in (11) reads as
@@
In the following, some assumptions on the optimization algorithm are provided. By denoting with $z_{\star}(x)$ the unique solution of the problem (10) at given initial conditions $x$, we assume as follows that $z_{\star}(x)$ is a globally exponentially stable equilibrium point of the optimizer dynamics (11).
@@
$z_{\star}(x)=\mathcal{T}(z_{\star}(x),x).$ (13)
Moreover, there exists a continuous function $\mathcal{L}:\mathbf{Z}\to\mathbb{R}_{+}$ such that
$b_{1}\left\|z-z_{\star}(x)\right\|^{2}\leq\mathcal{L}(z)\leq b_{2}\left\|z-z_{\star}(x)\right\|^{2}$ (14a)
$\mathcal{L}(\mathcal{T}(z,x))-\mathcal{L}(z)\leq-b_{3}\left\|z-z_{\star}(x)\right\|^{2}$ (14b)
$\mathcal{L}(z)-\mathcal{L}(z^{\prime})\leq b_{4}\left\|z-z^{\prime}\right\|(\left\|z-z_{\star}(x)\right\|+\left\|z^{\prime}-z_{\star}(x)\right\|),$ (14c)
for all $z,z^{\prime}\in\mathbf{Z}$, $x\in\mathbf{X}$, and some $b_{1},b_{2},b_{3},b_{4}>0$. Further, the functions $\mathcal{T}$ and $z_{\star}$ are $L_{\mathcal{T}}$- and $L_{z_{\star}}$-Lipschitz continuous, respectively, for some $L_{\mathcal{T}},L_{z_{\star}}\!>\!0$.
As is standard in MPC (see, e.g., [3]), we assume that the optimal solution to (10) yields an exponentially stable equilibrium $x_{\star}\in\mathbf{X}$ for the reduced system.
@@
There exists a continuous function $W:\mathbf{X}\to\mathbb{R}$ and $\bar{\delta}_{3}>0$ such that for all $\delta\in(0,\bar{\delta}_{3})$
@@
$W(f_{R}(x,\Pi(z_{\star}(x)),\delta))-W(x)\leq-\delta c_{3}\left\|x-x_{\star}\right\|^{2}$ (15b)
$W(x)-W(x^{\prime})\leq c_{4}\left\|x-x^{\prime}\right\|(\left\|x\right\|+\left\|x^{\prime}\right\|),$ (15c)
for all $x,x^{\prime}\in\mathbf{X}$ and some $c_{1},c_{2},c_{3},c_{4}>0$.
@@ III-B Closed-Loop System: Stability Guarantees
The proof of Theorem 1 is provided in Section IV-C after the preparatory analysis developed in Section IV.
@@ IV THEORETICAL ANALYSIS
In this section, we analyze the closed-loop system arising from the interconnection of the plant dynamics (1) and the optimizer one (11) to prove Theorem 1. More in detail, we interpret this closed-loop system as a two-time-scale system in which the fast subsystem includes both the plant extra dynamics (1b) and the optimizer dynamics (11), while the slow part includes only the plant target dynamics (1a). This feedback interconnection is graphically depicted in Fig. 2.
@@
Hence, as customary in two-time-scale system analysis, we study (i) the fast subsystem in the ideal case in which the slow state $x_{{t}}$ is fixed (cf. Section IV-B) and, conversely, (ii) the slow one in the ideal case in which the fast state $(\xi_{{t}},z_{{t}})$ lies in its equilibrium manifold (cf. Section IV-B). Finally, in Section IV-C we combine the results obtained in this preparatory phase to prove Theorem 1.
@@ IV-A Boundary-Layer System Analysis
$\tilde{\xi}_{{t}+1}$ $=g(\tilde{\xi}_{t}+\xi_{\text{eq}}(x,\Pi(\tilde{z}_{t}+z_{\star}(x))),x,\Pi(\tilde{z}_{t}+z_{\star}(x)),\delta)$
$\quad-\xi_{\text{eq}}(x,\Pi(\mathcal{T}(\tilde{z}_{t}+z_{\star}(x),x)))$ (17a)
$\tilde{z}_{{t}+1}$ $=\mathcal{T}(\tilde{z}_{t}+z_{\star}(x),x)-z_{\star}(x).$ (17b)
For the sake of compactness, let $\tilde{g}:\boldsymbol{\Xi}\times\mathbf{X}\times\mathbf{Z}\times\mathbb{R}_{+}\to\boldsymbol{\Xi}$, $\tilde{\mathcal{T}}:\mathbf{Z}\times\mathbf{X}\to\mathbf{Z}$, and $\Delta\xi_{\text{eq}}:\mathbf{X}\times\mathbf{Z}\to\boldsymbol{\Xi}$ be defined as
$\tilde{g}(\tilde{\xi},x,\Pi(\tilde{z}),\delta)$ $:=g(\tilde{\xi}+\xi_{\text{eq}}(x,\Pi(\tilde{z})),x,\Pi(\tilde{z}),\delta)$
$)\displaystyle\hskip 11.38092pt-\xi_{\text{eq}}(x,\Pi(\tilde{z}$)) (18a)
$\tilde{\mathcal{T}}(\tilde{z},x)$ $:=\mathcal{T}(\tilde{z}+z_{\star}(x),x)-z_{\star}(x)$ (18b)
@@
$\tilde{\xi}_{{t}+1}$ $=\tilde{g}(\tilde{\xi}_{t},x,\Pi(\tilde{z}_{t}+z_{\star}(x)),\delta)\!+\!\Delta\xi_{\text{eq}}(x,\tilde{z}_{{t}+1},\tilde{z}_{t})$ (19a)
$\tilde{z}_{{t}+1}$ $=\tilde{\mathcal{T}}(\tilde{z}_{t},x).$ (19b)
The following lemma ensures global exponential stability of the origin for the boundary-layer system (19).
@@
There exists a continuous function $\mathcal{U}:\boldsymbol{\Xi}\times\mathbf{Z}\to\mathbb{R}$ such that, for all $\delta\in(0,\min\{\bar{\delta}_{1},\bar{\delta}_{2}\})$, it holds
$d_{1}(\|\tilde{\xi}\|^{2}+\|\tilde{z}\|^{2})$ $\leq\mathcal{U}(\tilde{\xi},\tilde{z})\leq d_{2}(\|\tilde{\xi}\|^{2}+\|\tilde{z}\|^{2})$ (20a)
@@
$),\displaystyle\quad\times(\|\operatorname{\mathrm{col}}(\tilde{\xi},\tilde{z})\|+\|\operatorname{\mathrm{col}}(\tilde{\xi}^{\prime},\tilde{z}^{\prime}$$\|)$,
for all $\tilde{\xi}\in\boldsymbol{\Xi}$, $\tilde{z}\in\mathbf{Z}$, $x\in\mathbf{X}$, and some constants $d_{1},d_{2},d_{3},d_{4}>0$, where $\tilde{\xi}_{+}:=\tilde{g}(\tilde{\xi},x,\Pi(\tilde{z}+z_{\star}(x)),\delta)+\Delta\xi_{\text{eq}}(x,\tilde{z},\tilde{\mathcal{T}}(\tilde{z},x))$ and $\tilde{z}_{+}:=\tilde{\mathcal{T}}(\tilde{z},x)$.
The proof of Lemma 1 is provided in Appendix VI-B.

### v2

@@ I INTRODUCTION
The main drawback of MPC schemes lies in their computational complexity, which stems from the need to solve an optimal control problem online at each sampling instant. Consequently, a central challenge in current MPC research is to mitigate this limitation. In this context, we focus on two strategies that are widely adopted in both theoretical studies and practical implementations: (i) model order reduction, and (ii) suboptimal control approaches. When dealing with complex systems characterized by hierarchical structures or multiple interconnected subsystems, the computational complexity of MPC can become prohibitive. In such cases, reduced-order models can be employed to simplify the optimization problem while still capturing the essential dynamics of the system. In this spirit, works [4, 5, 6] propose multi-rate control architectures, in which high-level routines based on MPC operate on reduced-order models and provide setpoints to low-level controllers acting on the full-order ones. This paradigm, which is very common in robotic applications, has been formalized in the tutorial work [7]. In [8], the authors guarantee the stability of the full closed-loop system despite the use of a reduced model in the MPC optimizer. A reduced-order MPC scheme is presented in [9], where stability and constraint satisfaction are proven while explicitly accounting for model reduction errors, whose bounds can be computed, e.g., as in [10]. In [11], Galerkin projection methods are employed to operate on a low-dimensional surrogate model. The authors of [12] present a robust MPC framework for two-time-scale linear systems, where only a reduced-order model neglecting the fast dynamics is used in the MPC optimizer.
@@ II SCENARIO DESCRIPTION
where $x_{{t}}\in\mathbf{X}\subseteq\mathbb{R}^{n}$ is the target state, $\xi_{{t}}\in\boldsymbol{\Xi}\subseteq\mathbb{R}^{p}$ is the extra dynamics state, $f:\mathbf{X}\times\boldsymbol{\Xi}\times\mathbf{U}\times\mathbb{R}_{+}\to\mathbf{X}$ and $g:\boldsymbol{\Xi}\times\mathbf{X}\times\mathbf{U}\times\mathbb{R}_{+}\to\boldsymbol{\Xi}$ denote the target and extra dynamics, respectively, $u_{{t}}\in\mathbf{U}\subseteq\mathbb{R}^{m}$ is the control input, and $\delta>0$ is a parameter that allows to arbitrarily modulate the relative speed between the target state $x_{{t}}$ and the extra dynamics state $\xi_{{t}}$. We characterize this property as follows.
@@
There exist $\bar{\delta}_{1},L_{f}>0$ such that, for all $\delta\in(0,\bar{\delta}_{1})$, it holds
@@
$\left\|g(\xi,x,u,\delta)-g(\xi^{\prime},x,u^{\prime},\delta)\right\|$ $\leq L_{g}\left\|\xi\!-\!\xi^{\prime}\right\|+L_{g}\left\|u\!-\!u^{\prime}\right\|,$
for all $\xi,\xi^{\prime}\in\boldsymbol{\Xi}$, $x\in\mathbf{X}$, $u,u^{\prime}\in\mathbf{U}$.
@@
$\!a_{1}\!\|\xi\!-\!\xi_{\text{eq}}(x,u)\|^{2}\!\!\leq\!\mathcal{G}(\xi\!-\!\xi_{\text{eq}}(x,u)\!)\!\leq\!a_{2}\|\xi\!-\!\xi_{\text{eq}}(x,u)\|^{2}\!\!$ (3a)
$\mathcal{G}\big(g(\xi,x,u,\delta)-\xi_{\text{eq}}(x,u)\big)-\mathcal{G}(\xi-\xi_{\text{eq}}(x,u))\leq-a_{3}\|\xi-\xi_{\text{eq}}(x,u)\|^{2},$ (3b)
for all $\xi\in\boldsymbol{\Xi}$, $x\in\mathbf{X}$, $u\in\mathbf{U}$ and some constants $a_{1},a_{2},a_{3},a_{4}>0$.
@@
By adopting a Forward-Euler discretization with sampling time $\delta>0$ and scaling the armature inductance as $L=\tilde{L}\delta$ with $\tilde{L}>0$, we get a discretized counterpart of (5) given by
@@
Indeed, by observing (7), we can see that system (6) satisfies Assumption 1 with $L_{f}=K_{t}/J$ and $L_{g}=\max\{|1-R/\tilde{L}|,K_{e}/\tilde{L},1/\tilde{L}\}$. Moreover, by observing (7b), we also note that subsystem (6b) admits an equilibrium function $\xi_{\text{eq}}:\mathbb{S}^{1}\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ defined as
@@
Remark 1
We emphasize the dual role of the parameter $\delta$ in our framework. In classical continuous-time singular perturbation theory (see, e.g., [22]), timescale separation is typically induced by a small physical parameter that accelerates the fast subsystem. In the proposed discrete-time setting, this role is naturally played by the sampling time $\delta$. Indeed, as illustrated by the discretization process in Example 1, choosing a sufficiently small sampling time induces, from a mathematical viewpoint, a timescale separation between the fast extra dynamics and the slow target one, provided that the speed of the former can be arbitrarily tuned with respect to $\delta$.
III SUBOPTIMAL AND REDUCED-ORDER MPC: STRATEGY DESCRIPTION AND
STABILITY GUARANTEES
@@ III-A MPC Strategy Description
The control input $u_{{t}}$ is iteratively computed through a suboptimal MPC scheme applied to the reduced system, thereby adhering to and further reinforcing the rationale of computational efficiency motivating the use of $f_{R}$. Specifically, given a time horizon $T\in\mathbb{N}$, at each iteration ${t}\in\mathbb{N}$, we consider the finite-horizon optimal control problem
$\begin{aligned} \min_{\{u_{{\tau}}\}_{{\tau}=0}^{T-1}}\quad&\sum_{{\tau}=0}^{T-1}\ell\left(x_{{\tau}},u_{{\tau}}\right)+\ell_{T}\left(x_{T}\right)\\ \text{s.t.}\quad&x_{0}=x_{{t}}\\ &x_{{{\tau}+1}}=f_{R}(x_{{\tau}},u_{{\tau}},\delta_{\text{MPC}}),\quad{\tau}=0,\ldots,T-1\\ &u_{{\tau}}\in\mathcal{U},\quad x_{{{\tau}+1}}\in\mathcal{X},\quad{\tau}=0,\ldots,T-1\\ &x_{T}\in\mathcal{X}_{f},\end{aligned}$ (10)
where $\ell:\mathbf{X}\times\mathbf{U}\to\mathbb{R}_{+}$ is the stage cost, $\ell_{T}:\mathbf{X}\to\mathbb{R}_{+}$ is the terminal cost, $\mathcal{U}\subseteq\mathbf{U}$ is the input constraint set, $\mathcal{X}\subseteq\mathbf{X}$ is the state constraint set, $\mathcal{X}_{f}\subseteq\mathbf{X}$ is the terminal constraint set, and $\delta_{\text{MPC}}>0$ is the sampling time used by the optimizer. Problem (10) is addressed suboptimally. More in detail, we consider an optimizer dynamics of the form
$z_{{{t}+1}}$ $=\mathcal{A}(z_{{t}},x_{{t}})$ (11a)
@@
where $z_{{t}}\in\mathbf{Z}\subseteq\mathbb{R}^{Tm}$ is the optimizer state, $\mathcal{A}:\mathbf{Z}\times\mathbf{X}\to\mathbf{Z}$ describes the optimization algorithm dynamics, and $T\in\mathbb{N}$ is the prediction horizon. As customary in MPC schemes, only the first element of the $z_{{t}}$ sequence is applied to the plant as control input, hence the projection map $\Pi:\mathbf{Z}\to\mathbf{U}$ in (11) reads as
@@
In the following, some assumptions on the optimization algorithm are provided. By denoting with $z_{\star}(x)$ the unique solution to problem (10) at given initial conditions $x$, we assume as follows that $z_{\star}(x)$ is a globally exponentially stable equilibrium point of the optimizer dynamics (11).
@@
$z_{\star}(x)=\mathcal{A}(z_{\star}(x),x).$ (13)
Moreover, there exist a continuously differentiable function $\mathcal{L}:\mathbb{R}^{Tm}\to\mathbb{R}_{+}$ with $b_{4}$-Lipschitz continuous gradient $\nabla_{z}\mathcal{L}(z)$ and constants $b_{1},b_{2},b_{3},b_{4}>0$ such that
$b_{1}\left\|z\!\!-\!z_{\star}(x)\right\|^{2}\!\leq\!\mathcal{L}(z\!\!-\!z_{\star}(x))\leq b_{2}\left\|z\!\!-\!z_{\star}(x)\right\|^{2}$ (14a)
$\mathcal{L}(\mathcal{A}(z,x)\!\!-\!z_{\star}(x))\!-\!\mathcal{L}(z\!\!-\!z_{\star}(x))\!\leq\!-b_{3}\left\|z\!\!-\!z_{\star}(x)\right\|^{2}\!,\!$ (14b)
for all $z\in\mathbf{Z}$ and $x\in\mathbf{X}$.
Furthermore, we assume that the optimal solution of (10) yields an exponentially stable equilibrium $x_{\star}\in\mathbf{X}$ for the reduced system.
@@
There exist a continuously differentiable function $W:\mathbf{X}\to\mathbb{R}$ with $c_{4}$-Lipschitz continuous gradient $\nabla_{x}W(x)$, constants $c_{1},c_{2},c_{3}>0$, and $\bar{\delta}_{3}>0$ such that, for all $\delta\in(0,\bar{\delta}_{3})$, it holds
@@
$W(f_{R}(x,\Pi(z_{\star}(x)),\delta))\!-\!W(x)\!\leq\!-\delta c_{3}\left\|x-x_{\star}\right\|^{2}\!,\!$ (15b)
for all $x\in\mathbf{X}$. Moreover, for all $\delta\in(0,\bar{\delta}_{3})$ and some $L_{f_{R}}>0$, it holds
$\left\|f_{R}(x,\Pi(z_{\star}(x)),\delta)-x\right\|$ $\leq\delta L_{f_{R}}\left\|x-x_{\star}\right\|,$ (16)
for all $x\in\mathbf{X}$.
Remark 2
Sufficient conditions on the underlying optimal control problem (10) can be established to guarantee Assumption 4, namely, exponential stability of $x_{\star}$ for the reduced-order dynamics $x_{{{t}+1}}=f_{R}(x_{{t}},u_{{t}},\delta)$ under the optimal control law $u_{{t}}=\Pi(z_{\star}(x_{{t}}))$. These conditions regard, e.g., appropriate choice of stage and terminal cost functions, terminal constraints, sufficiently long prediction horizon, and regularity conditions on the system dynamics, see, e.g., [3, Chap. 2, pp. 139–141].
@@ III-B Closed-Loop System: Stability Guarantees
The proof of Theorem 1 will be provided in a forthcoming document. A sketch of the proof is provided in Section IV-C.
@@ IV THEORETICAL ANALYSIS
In this section, we analyze the closed-loop system arising from the interconnection of the plant dynamics (1) and the optimizer one (11) to prove Theorem 1. Specifically, we interpret this closed-loop system as a two-time-scale system in which the fast subsystem includes both the plant extra dynamics (1b) and the optimizer dynamics (11), while the slow part includes only the plant target dynamics (1a). This feedback interconnection is graphically depicted in Fig. 2.
@@
Hence, as customary in two-time-scale systems analysis, we study (i) the fast subsystem in the ideal case in which the slow state $x_{{t}}$ is fixed (cf. Section IV-A) and, conversely, (ii) the slow one in the ideal case in which the fast state $(\xi_{{t}},z_{{t}})$ lies in its equilibrium manifold (cf. Section IV-B). Finally, in Section IV-C we combine the results obtained in this preparatory phase to prove Theorem 1.
@@ IV-A Boundary-Layer System Analysis
$\tilde{\xi}_{{t}+1}$ $=g(\tilde{\xi}_{t}+\xi_{\text{eq}}(x,\Pi(\tilde{z}_{t}+z_{\star}(x))),x,\Pi(\tilde{z}_{t}+z_{\star}(x)),\delta)-\xi_{\text{eq}}(x,\Pi(\mathcal{A}(\tilde{z}_{t}+z_{\star}(x),x)))$ (18a)
$\tilde{z}_{{t}+1}$ $=\mathcal{A}(\tilde{z}_{t}+z_{\star}(x),x)-z_{\star}(x).$ (18b)
For the sake of compactness, we define the functions
$\tilde{g}(\tilde{\xi},x,\tilde{z},\delta)$ $:=g(\tilde{\xi}+\xi_{\text{eq}}(x,\Pi(\tilde{z}+z_{\star}(x))),x,\Pi(\tilde{z}+z_{\star}(x)),\delta)-\xi_{\text{eq}}(x,\Pi(\tilde{z}+z_{\star}(x)))$ (19a)
$\tilde{\mathcal{A}}(\tilde{z},x)$ $:=\mathcal{A}(\tilde{z}+z_{\star}(x),x)-z_{\star}(x)$ (19b)
@@
$\tilde{\xi}_{{t}+1}$ $=\tilde{g}(\tilde{\xi}_{t},x,\tilde{z}_{t},\delta)\!+\!\Delta\xi_{\text{eq}}(x,\tilde{z}_{{t}+1},\tilde{z}_{t})$ (20a)
$\tilde{z}_{{t}+1}$ $=\tilde{\mathcal{A}}(\tilde{z}_{t},x).$ (20b)
The following lemma ensures global exponential stability of the origin for the boundary-layer system (20). To properly state this result, let us introduce the set $\mathcal{S}$ defined as
$\mathcal{S}:=$ $\{(\tilde{\xi},x,\tilde{z})\in\mathbb{R}^{p}\times\mathbf{X}\times\mathbb{R}^{Tm}\mid\tilde{\xi}+\xi_{\text{eq}}(x,\Pi(\tilde{z}+z_{\star}(x)))\in\boldsymbol{\Xi},\tilde{z}+z_{\star}(x)\in\mathbf{Z}\}.$
@@
There exists a continuous function $U:\mathbb{R}^{p}\times\mathbb{R}^{Tm}\to\mathbb{R}$ such that, for all $\delta\in(0,\min\{\bar{\delta}_{1},\bar{\delta}_{2}\})$, it holds
$d_{1}(\|\tilde{\xi}\|^{2}+\|\tilde{z}\|^{2})$ $\leq U(\tilde{\xi},\tilde{z})\leq d_{2}(\|\tilde{\xi}\|^{2}+\|\tilde{z}\|^{2})$ (21a)
@@
for all $(\tilde{\xi},x,\tilde{z})\in\mathcal{S}$, and some constants $d_{1},d_{2},d_{3},d_{4}>0$, where $\tilde{\xi}_{+}:=\tilde{g}(\tilde{\xi},x,\tilde{z},\delta)+\Delta\xi_{\text{eq}}(x,\tilde{z},\tilde{\mathcal{A}}(\tilde{z},x))$ and $\tilde{z}_{+}:=\tilde{\mathcal{A}}(\tilde{z},x)$.
The proof of Lemma 1 will be provided in a forthcoming document.